You are on planet Dither, whose inhabitants often change their minds on how to choose a reference axis. At time t= 0, while standing 3.0 m to the left of the origin of a reference axis for which the positive direction points to the left, you launch a toy car that moves with increasing speed to the right. The car's x

x coordinate (x component of the position) is given by x(t)=p+qt+rt

2

x(t)=p+qt+rt2, where p

p= +3.0m

m, q

qq= -2.0m/s

m/s, and r

rr= -2.0m/s

2

m/s2.


What is the equation of the car's x coordinate if the reference axis is flipped so that the positive direction points to the right?

Express your answer in terms of some or all of the variables p q t r



1
Expert's answer
2021-03-26T08:03:56-0400

"x(t)=p+qt+rt^2,"

"x(t)=3-2t-2t^2,"


"x'(t)=-x(t)=-p-qt-rt^2,"

"x'(t)=2t^2+2t-3."